The path that a motorcross dirt bike rider follows during a jump is given by y=-0.4x^2+4x+10 where x is the horizontal distance in feet from the edge of the ramp and y is the height in feet.
a.Graph the function on a coordinate plane
b.What is the maximum height of the rider during the jump? Explain how you know.

Answer by Alan3354(69443)

The path that a motorcross dirt bike rider follows during a jump is given by y=-0.4x^2+4x+10 where x is the horizontal distance in feet from the edge of the ramp and y is the height in feet.
a.Graph the function on a coordinate plane
--
Pick values for x and find y.
Plot the points.
Draw a curve thru them.
=================
b.What is the maximum height of the rider during the jump?
It's the vertex of the parabola at x = -b/2a = -4/-0.8 = 5
y=-0.4x^2+4x+10 = -0.4*25 + 4*5 + 10 = 20 feet
-----
Explain how you know.
How I know what?